OpenStudy (anonymous):

I don't know how to calculate the rotational kinetic energy. I = 524.3 and omega = 6 rad/sec

OpenStudy (anonymous):

Hmm.. same as the translational kinetic energy is\[E_{kin} = \frac{m \, v^2}{2}\] the rotational kinetic energy is \[E_{kin} = \frac{I \, \omega^2}{2}\]

OpenStudy (anonymous):

for omega, do i square just the 6 or (2* pi * 6)

OpenStudy (anonymous):

Why would you multiply it with 2π? ω is your angular velocity... You are thinking of \[\omega = 2 \pi f\]This identity is very useful and it's very good, you know it, but you are already given ω in this case. Note also, that the unit of f is Hz and the unit of ω is rad/s. While both are technically just 1/s, we use the different unit-names to make the distinction even easier.